You want a seat on the board of directors of Zeph, Inc. The company has 305,000 shares of stock outstanding and the stock sells for $52 per share. There are currently 4 seats up for election. If the company uses cumulative voting, how much will it cost you to guarantee that you will be elected to the board?
Answer #1

Percent of stock needed = 1 / (N + 1)

Percent of stock needed = 1 / (4 + 1)

Percent of stock needed = 0.20, or 20%

Number of shares to purchase = (305,000 × 0.20) + 1

Number of shares to purchase = 61,001

Total cost = 61,001 × $52

Total cost = $3,172,052